OpenStudy (anonymous):

find the slope-intercept equation, for the line with the indicated slope and y-intercept. slope 5 over 4 y-intercept(0,-7)

OpenStudy (anonymous):

y - (-7) = -5/4 ( x - 0 ) y + 7 = -5/4x y = -5/4x - 7

OpenStudy (anonymous):

I thought when you write 5/4 that means multiply?

OpenStudy (anonymous):

thats divide or a fraction

OpenStudy (anonymous):

o yea ok thanks